OpenStudy (goformit100):

Does there exist a function which is continuous everywhere but not differentiable at exactly two points? Justify your answer .

OpenStudy (goformit100):

@ajprincess @lalaly

OpenStudy (goformit100):

@elizabeth080808

OpenStudy (anonymous):

\[y=\left| \sin(x) \right|\],\[0\le x \le 3\pi\]

OpenStudy (goformit100):

Thankyou Sir

OpenStudy (anonymous):

y is continuous everywhere but not differentiable at \[x=\pi\] and \[x=2\pi\] as there exists a sharp edge at these points

OpenStudy (goformit100):

ok

OpenStudy (anonymous):

and i am a girl not sir......